OpenStudy (anonymous):

Find the standard form of the equation of the ellipse and find the center. Given the vertices:(0,+-8); Foci: (0,+-4)

OpenStudy (mertsj):

If the vertices are (0,8) and (0,-8), where is the center?

OpenStudy (anonymous):

(0,0)

OpenStudy (mertsj):

Exactly.

OpenStudy (mertsj):

And what is the value of "a"

OpenStudy (anonymous):

8?

OpenStudy (mertsj):

yes. And do you know that the distance from the center to the focus is "c" ?

OpenStudy (anonymous):

yes

OpenStudy (mertsj):

So "c" is ???

OpenStudy (anonymous):

so you do c^2=a^2-b^2?

OpenStudy (mertsj):

yes

OpenStudy (mertsj):

And remember that the major axis is vertical so a^2 goes under y^2

OpenStudy (anonymous):

but since the foci is given...c=-+2

OpenStudy (mertsj):

The focal point is (0,4) so c = 4

OpenStudy (anonymous):

oh you square the 2

OpenStudy (mertsj):

Where did you get 2?

OpenStudy (anonymous):

so what goes under x^2?

OpenStudy (mertsj):

c is the distance from the center to the focal point. If the center is (0,0) and the focal point is (0,4), then c has to be 4

OpenStudy (anonymous):

oh ok

OpenStudy (mertsj):

16=64-b^2 -48=-b^2 48=b^2

OpenStudy (anonymous):

major axis is the same as vertices?

OpenStudy (mertsj):

The major axis is the distance between the vertices and its length is 2a

OpenStudy (anonymous):

so b=sqrt48

OpenStudy (mertsj):

Yes.